OpenStudy (anonymous):

The are of quadrilateral formed by \[5x \pm2y \pm5=0\]

OpenStudy (anonymous):

I got 5 square units

OpenStudy (anonymous):

u r correct

OpenStudy (anonymous):

Can u shw ???

OpenStudy (anonymous):

These are the coordinate points of the given quadrilateral 

OpenStudy (anonymous):

Using application of determinant for area we can find out the answer

OpenStudy (anonymous):

Lol...we will be having four diff eq

OpenStudy (anonymous):

Ok here is it \[\frac{ 1 }{ 2 }|\left[\begin{matrix}0 & \frac{ 5 }{ 2 } \\ 1 & 0\end{matrix}\right]+\left[\begin{matrix}1 & 0 \\ 0 & \frac{ -5 }{ 2 }\end{matrix}\right]+\left[\begin{matrix}0 & \frac{ -5 }{ 2 }\ \\ -1 & 0\end{matrix}\right]+\left[\begin{matrix}-1 & 0 \\ 0 & \frac{ 5 }{ 2 }\end{matrix}\right]|\]

OpenStudy (anonymous):

thxx

OpenStudy (anonymous):

@Zekarias sorry But I didnt get what u did?

OpenStudy (anonymous):

That was the application of determinant